OpenStudy (anonymous):

Will Medal And Fan ! Question 8.8. Select the pair of equations whose graphs are perpendicular.(Points : 5) A. 2x – 8y = 9 12x – 3y = 7 B. 3x + 6y = 8 y = 2x – 8 C. y = 2x – 7 x + y = 3 D. 2y = –3x + 5 2x + 3y = 4

OpenStudy (nerdgirl):

The answer is D.

OpenStudy (anonymous):

D

OpenStudy (anonymous):

thanks guys <3

OpenStudy (anonymous):

Could you fan me.

OpenStudy (anonymous):

already did dude (:

OpenStudy (anonymous):

Oh. Thx! :D

OpenStudy (anonymous):

anytime !

OpenStudy (dan815):

the answer is not D

OpenStudy (dan815):

2 lines are perpendicular when their slopes are negative reciprocals of each other

OpenStudy (dan815):

B. 3x + 6y = 8 y = 2x – 8 if u arrange 3x+6y=8 in terms of y y=-3x/6 +8/6 =-1/2 x +4/3, y=2x-8 and -1/2 is the NEGATIVE RECIPROCAL OF 2/1 so

OpenStudy (anonymous):

oh. what is it? i got d on a line graph.

OpenStudy (dan815):

y=(-1/2) x +4/3, y=(2/1)x-8 ^--------------^------------- slopes

OpenStudy (anonymous):

wait what? I'm confused :/

OpenStudy (anonymous):

@dan815

OpenStudy (dan815):

okay i can explain

OpenStudy (dan815):

but do u have time for that, arent u writing an exam

OpenStudy (anonymous):

I have 21 minutes left and only one question so i have time

OpenStudy (dan815):

okay well

OpenStudy (dan815):

you know about line form? y=mx+b m is the slope b is the y intercept

OpenStudy (anonymous):

Yes I understand that

OpenStudy (dan815):

okay so lets see what this m really means

OpenStudy (dan815):

lets look at y=mx line this m is slope means change in y / change in x in other words you go up by m units for every 1 unit of x

OpenStudy (anonymous):

I just don't get how D isn't the answer

OpenStudy (dan815):

because when you rearrange it in y=mx+b form the slopes are not negative recriprocals of each other

OpenStudy (dan815):

D. 2y = –3x + 5, 2x + 3y = 4 y=-3/2x+5/2, y=-2x/3 +4/3 ^---------------------^--------- it has to be -3/2 and 2/3 or 3/2 and -2/3

OpenStudy (dan815):

but in this case both of the slope are negative

OpenStudy (dan815):



OpenStudy (dan815):

ud expect a line perpendicular to be something like htis 

OpenStudy (anonymous):

So the answer would be B then, right?

OpenStudy (dan815):

yeah.. but -.-

OpenStudy (dan815):

do u really understand

OpenStudy (anonymous):

I do. You've helped a lot, thank you! <3